Find the area of an oak cabinet that is a large square with sides of 6x feet and has a square cut-out region whose sides are y feet. Write your answers in factored form.
I think the equation is 6x⋅6x-y⋅y or 6x^2-y^2, but I'm not entirely sure if that is right or how to put it in factored form.

Thank you in advance for your help!

Answer by Theo(13342)

i think the area of the large square is (6x)^2 = 36x^2

i think the area of the small square that is cut out of it is y^2

the net area of the cabinet would be 36x^2 - y^2.

this takes the form of a^2 - b^2 = (a-b) * (a+b), which is a general formula that you might be familiar with.

a is equal to 6x.
b is equal to y

36x^2 - y^4 becomes (6x - y) * (6x + y)

to prove that this is true, simply multiply the factors together.

you get:

6x * 6x = 36x^2
6x * y = 6xy
-y * 6x = -6xy
-y * y = -y^2

the expression becomes 36x^2 + 6xy - 6xy - y^2

combine like terms to get 36x^2 - y^2.

the positive and negative 6xy cancel each other out and disappear.